# Scaling Efficient LLMs

### v4

@@ Abstract
Trained LLMs in the transformer architecture are typically sparse in that most of the parameters are negligible, raising questions on efficiency. Furthermore, the so called "AI scaling law" for transformers suggests that the number of parameters must scale linearly with the size of the data. In response, we inquire into efficient LLMs, i.e. those with the fewest parameters that achieve the desired accuracy on a training corpus. Specifically, by comparing theoretical and empirical estimates of the Kullback-Liebler divergence, we derive a natural AI scaling law that the number of parameters in an efficient LLM scales as $D^{\gamma}$ where $D$ is the size of the training data and $\gamma\in[0.44,0.72]$, suggesting the existence of more efficient architectures. Against this backdrop, we propose recurrent transformers, combining the efficacy of transformers with the efficiency of recurrent networks, progressively applying a single transformer layer to a fixed-width sliding window across the input sequence. Recurrent transformers (a) run in linear time in the sequence length, (b) are memory-efficient and amenable to parallel processing in large batches, (c) learn to forget history for language tasks, or accumulate history for long range tasks like copy and selective copy, and (d) are amenable to curriculum training to overcome vanishing gradients. In our experiments, we find that recurrent transformers perform favorably on benchmark tests.
@@ 1 Introduction
LLMs are artificial neural networks typically built on the transformer architecture as proposed by Vaswani et al. (2017). Recent LLMs have hundreds of billions of parameters, e.g.,Brown et al. (2020), Rae et al. (2021), Smith et al. (2022), Thoppilan et al. (2022). However, trained transformer LLMs are sparse in that most of the parameters are negligible, Gent (2023), raising the question as to whether the models really need to be large to perform, thereby consuming vast computational resources. Furthermore, the so called "AI scaling law" for transformers suggests that the number of parameters must scale linearly with the size of the data, and that we are running out of usable training data to build even larger models, e.g., Kaplan et al. (2020), Hoffmann et al. (2022), Muennighoff et al. (2023). In response, we inquire into efficient LLMs independent of architecture, i.e. those with the fewest parameters that achieve the desired accuracy on a training corpus.
@@
We compare theoretical and empirical estimates of the Kullback-Liebler divergence, Kullback & Leibler (1951), to show that at prevalent sequence lengths, the number of unique sequences in a natural training corpus of size $D$ is between $o(D^{0.44})$ and $\mathcal{O}(D^{0.72})$. Our result leverages the empirical estimates of Hoffmann et al. (2022). From a learning theoretic perspective, the number of unique sequences as a function of the size of a training corpus is an upper bound on the dimension of the space of functions required to learn any natural data set of that size. Theoretically, the number of parameters in an efficient LLM would scale linearly with the dimension of the space, and hence linearly with the number of unique sequences. In contrast to the so-called AI scaling law for the transformer architecture that suggests the number of parameters must scale linearly with the size of the data, a salient implication of our result is that the number of parameters in an efficient LLM scales as $D^{\gamma}$ for $0.44\leq\gamma\leq 0.72$, suggesting the existence of more efficient LLM architectures.
@@
Our result also has implications for so-called emergent abilities i.e. “abilities that are not present in smaller-scale models but are present in large-scale models; thus they cannot be predicted by simply extrapolating the performance improvements on smaller-scale models” e.g., Wei et al. (2022), Ganguli et al. (2022), Arora & Goyal (2023) and Srivastava et al. (2023). Conversely, Schaeffer et al. (2023) find that LLMs may exhibit abilities that are latent but not new when the number of parameters is increased. Bridging the above, our result suggests that if the number of parameters of an LLM is smaller than the number of unique sequences in the training corpus, scaling up can uncover latent abilities.
@@
With our theoretical results as backdrop, we seek more efficient LLM architectures. Preceding transformers, Recurrent Neural Networks (RNNs), e.g., Elman (1990), their variants Long Short-Term Memory (LSTM), e.g., Hochreiter & Schmidhuber (1997) and Gated Recurrent Units (GRUs), e.g., Cho et al. (2014), were popular for linear time processing of sequential data streams such as natural language processing. More recent architectural alternatives include State Space Models, e.g., Gu et al. (2021), and minimal RNNs, Feng et al. (2024). Building on prior work, we propose recurrent transformers, combining the efficacy of transformers with the efficiency of recurrent networks, progressively applying a single transformer layer to a fixed-width sliding window across the input sequence. Prior work on recurrent transformers, e.g., Hutchins et al. (2022), focuses on improving the long-range performance of multi-layer transformers by making some of the layers recurrent with large block sizes and many additional weights. In contrast, our focus is to find small and efficient networks for a given task, in the form of a single recurrent layer with a scalar parameter that learns to accumulate or forget history. Specifically, our recurrent transformers (a) run in linear time in the sequence length; (b) are memory-efficient and amenable to parallel processing in large batches (c) learn to forget history for language tasks, or accumulate history for long-range tasks like copy and selective copy; and (d) are amenable to curriculum training to overcome vanishing gradients, e.g.,Hochreiter (1991), Sodhani et al. (2020). In our experiments, we find that a single-layer recurrent transformer can match the performance of multi-layer transformers on benchmark tests, delivering comparable accuracy with fewer parameters and much lower computational cost. Furthermore, the trainable accumulation capability enables a recurrent transformer to solve long range tasks at a fraction of the size of other recurrent architectures, e.g., Gu & Dao (2023), Ren et al. (2024), and at a fraction of the cost of regular transformers.
@@ 2 Background
An LLM is a statistical machine that takes as input a sequence of words of specified maximum length on a specified alphabet, and produces as output a probability distribution on possible next words. Training involves adjusting the model’s parameters while running through a body of non-duplicative text called the training corpus, to minimize the statistical error between the model and the corpus. The error in the model, also known as the loss function, is chosen as a continuous function to enable the use of stochastic gradient descent for adjusting the parameters of the model. The commonly used loss function involves the entropy of an infinite stream of words, as defined by Shannon (1948), Cover (1999). Intuitively, entropy is the measure of information in the stream and is a lower bound on the capacity of a channel required to transmit the stream.
@@
Let $S$ be the set of all sequences of length $l$ in the training corpus T. The Kullback-Liebler divergence of an LLM is
@@
The Kullback-Liebler divergence is well established to be non-negative. Rearranging, we get
@@
The term on the left above is the cross-entropy loss of the LLM on the training corpus. The first term on the right converges to the inherent entropy of the natural language on the training corpus. During the training process, the parameters of the LLM are iteratively adjusted to minimize the cross-entropy. In doing so, the LLM is effectively minimizing the Kullback-Liebler divergence, since the inherent entropy of the language is constant.
@@
We assume all calculations are performed in finite precision, where the unit interval is represented by the set $K=\{0,(1/k),(2/k),...,(k/k)\}$, for specified integer $k\geq 1$. To avoid overflow in loss calculations, division by 0 is replaced by division by $1/k.$
@@ 3 Empirical Scaling
Recent prior works report empirical expressions for the relationship between training loss and model scale, e.g., Kaplan et al. (2020)Hoffmann et al. (2022), Muennighoff et al. (2023).  Per current practice, e.g., Hoffmann et al. (2022), Xue et al. (2023), these works train LLMs on a single epoch of natural non-duplicative training corpora.  Specifically, Hoffmann et al. (2022) finds that for fixed sequence length, across a range of LLMs with $N$ parameters and training corpus of size $D$, the Kullback-Liebler divergence scales as follows
@@
Since each training run starts with a random initialization of the neural network, we assume Equation (3) holds with high confidence $(1-\delta)$, for small positive $\delta$ . The first term on the right reflects the ability of an LLM with $N$ parameters to capture the richness of the training corpus. The second term reflects the ability of the training corpus to capture the richness of the natural language.    Since $\alpha\approx\beta,A\approx B$, and the compute cost is $N\times D$, it follows that the optimal scaling strategy is $N\propto D$, which is also known as the AI Scaling Law.    However, $N\propto D$ suggests rote learning, raising the question of whether it is a natural requirement or a reflection of the transormer architecture used in the experiments.
@@ 4 Theoretical Scaling
We now consider LLMs in the context of the theoretical model of Probably Approximately Correct (PAC) learning, Valiant (1984). Recall that $S$ is the set of sequences of length $l$ in the training corpus $T$. Consider a hypothesis space of functions $F=\{f:S\rightarrow K\}$.
On a sequence $s$, let $f(s)$ represent the conditional probability distribution $Q(s)$ as in Equation 2. Define the discrete loss of $f$ as the probability that $f$ is incorrect on a sequence in $T$ with precision $k$, i.e.,
$L_{k}(f,T)=\sum_{p_{s}\neq q_{s}}p_{s}$ (4)
We follow Shalev-Shwartz & Ben-David (2014) in the definition below.
@@
Given $\epsilon,\delta\in(0,1)$, an agnostic learning algorithm finds $f\in F$ with confidence at least $(1-\delta)$such that
$L_{k}(f,T)\leq\min_{h\in F}L_{k}(h,T)+\epsilon$
There are several measures for the complexity of a space of functions that can be used to bound , such as the Vapnik Chervonenkis dimension, the Generalized dimension and Rademacher Complexity.  Since we need both upper and lower bounds on multi-class functions, we use the notion of the Generalized Dimension also known as the Natarajan dimension, Shalev-Shwartz & Ben-David (2014).   For brevity, we will simply refer to it as the dimension.
Definition 2.
(Generalized shattering) A set $C\subset X$ is shattered by $F$ if there exist two functions $f_{0},f_{1}$ in $F$ such that for every $x\in C$, $f_{0}(x)\neq f_{1}(x)$; and for every $B\subset C$, there exists $f_{2}\in F$, such that for all $x\in B,f_{2}(x)=f_{0}(x)$ , and for all $x\in(C-B)$, $f_{2}(x)=f_{1}(x)$.
Definition 3.
The dimension of a space of functions $F$ is the size of the largest set shattered by it, and is denoted by $dim(F)$.
Intuitively, the dimension of a space of functions is a measure of the richness of the space, i.e. the number of degrees of freedom across the functions in the space.
The following theorem using the asymptotic complexity notation of $\Theta$ is adapted from Shalev-Shwartz & Ben-David (2014) and proved therein.
@@
For fixed $k$ and $\delta$, the error in any agnostic learning algorithm for a space of functions $F$ from $m$ samples is  $\Theta[(dim(F)/m)^{0.5}]$
Therefore for fixed $k,\delta$, the expected test loss $L(f,P)$ of an agnostic learning algorithm is of the form
$L(f,P)=\min_{h\in F}L(h,P)+\Theta[(dim(F)/m)^{0.5}]$ (5)
Equation 5 is the theoretical analog to the empirical scaling law of Equation 3.   The first term on the right reflects the ability of a space of functions F to capture the richness of the training corpus. The second term reflects the ability of the training corpus to capture the richness of the natural language.
We now link the empirical bounds of Equation 3 with the theoretical bounds of Equation 5 to estimate the number of unique sequences in a training corpus as a function of its size. Since our results depend on the empirical estimates of Hoffmann et al. (2022), they hold in the context of the corresponding empirical conditions such as sequence length, corpus size and computational precision. Critically, our result is independent of the transformer architecture used to obtain the empirical estimates.
Consider an LLM with $N$ parameters and sequence length $l$ on a specific alphabet.   Recall that the empirical results of Hoffmann et al. (2022) are based on non-duplicative and natural training corpora and that $S$ is the set of sequences of length $l$ in the training corpus $T$.
@@
If Equation 3 holds for sequences of length $l$, the number of unique sequences $|S|$ of length $l$ in a corpus of size $D$ is between $o(D^{0.44})$ and $\mathcal{O}(D^{0.72})$.
@@
Consider an LLM with $N$ parameters representing a space of functions $F=\{f:S\rightarrow K\}$. Suppose the LLM computes a function $f$ after training on corpus $T$.  From Equation 2, we have
$\Delta=\sum_{s\in S}p_{s}log(p_{s}/q_{s})=\sum_{p_{s}\neq q_{s}}p_{s}log(p_{s}/q_{s})$
By Assumption 1, $log(p_{s}/q_{s})$ is maximum when $p_{s}=1,q_{s}=1/k$, and hence
$\Delta\leq\log(k)\sum_{p_{s}\neq q_{s}}p_{s}$
Substituting Equation 4, we have
$\Delta\leq\log(k)L_{k}(f,T)$
@@
$\Delta=\frac{A}{N^{\alpha}}+\frac{B}{D^{\beta}}\leq\log(k)L_{k}(f,T)$
Substituting Equation 5 in the above, we get the following where $d$ is the dimension of $F$.
$\frac{A}{N^{\alpha}}+\frac{B}{D^{\beta}}\leq\log(k)\left(\min_{h\in F}L_{k}(h,T)+\Theta[(d/D)^{0.5}]\right)$
As the number of parameters $N$ in the LLM goes to infinity, $A/N$ and $\min_{h\in F}L_{k}(h,T)$  go to zero. Furthermore, by definition, $d=dim(F)\leq|S|$. Taken together, we have
$\frac{B}{D^{\beta}}=\mathcal{O}\left(\frac{|S|}{D}\right)^{0.5}$
@@
$|S|=o(D^{1-2\beta})=o(D^{0.44}).$
Turning to the upper bound in the lemma, invoking Pinsker’s inequality, Csiszár & Körner (2011), we can write
$\Delta=\sum_{s\in S}p_{s}log(p_{s}/q_{s})\geq\frac{1}{2}\left(\sum_{s\in S}|p_{s}-q_{s}|\right)^{2}$
Substituting Equation 2 in the above, we get
$\Delta\geq\frac{1}{2}\left(\sum_{s\in S}\mathbb{P}\{\hat{s}\}|P(s)-Q(s)|\right)^{2}=\frac{1}{2}\left(\sum_{p_{s}\neq q_{s}}\mathbb{P}\{\hat{s}\}|P(s)-Q(s)|\right)^{2}$
Per Assumption 1, if $p_{s}\neq q_{s}$, $|P(s)-Q(s)|\geq 1/k$. Also, $P(s)\leq 1$. Substituting in the above,
$\Delta\geq\frac{1}{2k^{2}}\left(\sum_{p_{s}\neq q_{s}}\mathbb{P}\{\hat{s}\}\right)^{2}\geq\frac{1}{2k^{2}}\left(\sum_{p_{s}\neq q_{s}}\mathbb{P}\{\hat{s}\}P(s)\right)^{2}\geq\frac{1}{2k^{2}}\left(\sum_{p_{s}\neq q_{s}}p_{s}\right)^{2}$
Substituting Equation 4 in the above, we get
$\Delta\geq\frac{1}{2k^{2}}\left(L_{k}(f,T)\right)^{2}$
Substituting Equations 3 and 5 in the above,we get the following where $d$ is the dimension of $F$
$\frac{A}{N^{\alpha}}+\frac{B}{D^{\beta}}\geq\frac{1}{2k^{2}}\left(\min_{h\in F}L_{k}(h,T)+\Theta[(d/D)^{0.5}]\right)^{2}.$
As $N\rightarrow\infty$, $A/N$ and $\min_{h\in F}L_{k}(h,T)$ go to zero. Furthermore, by definition, $d=dim(F)\leq|S|$. Taken together, we get
$\frac{B}{D^{\beta}}=o\left(\frac{1}{2k^{2}}\frac{|S|}{D}\right)$
Substituting $\beta=0.28$ from Equation 3 in the above,  we get $|S|=\mathcal{O}(D^{1-\beta})=\mathcal{O}(D^{0.72}).$
From a learning theoretic perspective, the number of unique sequences as a function of the size of a training corpus is an upper bound on the dimension of the space of functions required to learn any natural dataset of that size. Theoretically, the number of parameters in an efficient LLM would scale linearly with the dimension of the space, and hence linearly with the number of unique sequences. The so-called AI scaling law for the transformer architecture suggests that the number of parameters must scale linearly with the size of the data, e.g., Kaplan et al. (2020), Hoffmann et al. (2022), Muennighoff et al. (2023). In contrast, our main result implies that the number of parameters of an efficient LLM need only be $\mathcal{O}(D^{\gamma})$ for $\gamma\in[0.44,0.72]$, suggesting the existence of substantially more efficient LLM architectures.
Our result also has implications for so-called emergent abilities i.e. “abilities that are not present in smaller-scale models but are present in large-scale models; thus they cannot be predicted by simply extrapolating the performance improvements on smaller-scale models” e.g., Wei et al. (2022), Ganguli et al. (2022), Arora & Goyal (2023) and Srivastava et al. (2023). Conversely, Schaeffer et al. (2023) find that LLMs may not exhibit new abilities simply by increasing the number of parameters. Our result implies that if the number of parameters of an LLM is smaller than the number of unique sequences in the training corpus, scaling up the LLM can uncover emergent skills.

### v5

@@ Abstract
Recent LLMs have hundreds of billions of parameters consuming vast resources. Furthermore, the so called "AI scaling law" for transformers suggests that the number of parameters must scale linearly with the size of the data. In response, we inquire into efficient LLMs, i.e. those with the fewest parameters that achieve the desired accuracy on a training corpus. Specifically, by comparing theoretical and empirical estimates of the Kullback-Leibler divergence, we derive a natural AI scaling law that the number of parameters in an efficient LLM scales as $D^{\gamma}$ where $D$ is the size of the training data and $\gamma\in[0.44,0.72]$, suggesting the existence of more efficient architectures. Against this backdrop, we propose recurrent transformers, combining the efficacy of transformers with the efficiency of recurrent networks, progressively applying a single transformer layer to a fixed-width sliding window across the input sequence. Recurrent transformers (a) run in linear time in the sequence length, (b) are memory-efficient and amenable to parallel processing in large batches, (c) learn to forget history for language tasks, or accumulate history for long range tasks like copy and selective copy, and (d) are amenable to curriculum training to overcome vanishing gradients. In our experiments, we find that recurrent transformers perform favorably on benchmark tests.
@@ 1 Introduction
LLMs are artificial neural networks typically built on the transformer architecture as proposed by VSPUJGK17. Recent LLMs have hundreds of billions of parameters consuming vast amount of resources, e.g.,bmrskdoa20, rbcmhsoi21, spnlrcoc22, tdhskcol22. Furthermore, the so called "AI scaling law" for transformers suggests that the number of parameters must scale linearly with the size of the data, and that we are running out of usable training data to build even larger models, e.g., kmhbccgrwa20, hbmbcros22, muennighoff2023scaling. In response, we inquire into efficient LLMs independent of architecture, i.e. those with the fewest parameters that achieve the desired accuracy on a training corpus.
@@
We compare theoretical and empirical estimates of the Kullback-Leibler divergence, kullback1951information, to show that at prevalent sequence lengths, the number of unique sequences in a natural training corpus of size $D$ is between $\Omega(D^{0.44})$ and $\mathcal{O}(D^{0.72})$. Our result leverages the empirical estimates of hbmbcros22. From a learning theoretic perspective, the number of unique sequences as a function of the size of a training corpus is an upper bound on the dimension of the space of functions required to learn any natural data set of that size. Theoretically, the number of parameters in an efficient LLM would scale linearly with the dimension of the space, and hence linearly with the number of unique sequences. In contrast to the so-called AI scaling law for the transformer architecture that suggests the number of parameters must scale linearly with the size of the data, a salient implication of our result is that the number of parameters in an efficient LLM scales as $D^{\gamma}$ for $0.44\leq\gamma\leq 0.72$, suggesting the existence of more efficient LLM architectures.
@@
Our result also has implications for so-called emergent abilities i.e. “abilities that are not present in smaller-scale models but are present in large-scale models; thus they cannot be predicted by simply extrapolating the performance improvements on smaller-scale models” e.g., wei2022emergent, ganguli2022predictability, arora2023theory and srivastava2023beyond. Conversely, schaeffer2023emergent find that LLMs may exhibit abilities that are latent but not new when the number of parameters is increased. Bridging the above, our result suggests that if the number of parameters of an LLM is smaller than the number of unique sequences in the training corpus, scaling up can uncover latent abilities.
@@
With our theoretical results as backdrop, we seek more efficient LLM architectures. Preceding transformers, Recurrent Neural Networks (RNNs), e.g., elman1990finding, their variants Long Short-Term Memory (LSTM), e.g., hochreiter1997long and Gated Recurrent Units (GRUs), e.g., cho2014learning, were popular for linear time processing of sequential data streams such as natural language processing. More recent architectural alternatives include State Space Models, e.g., gu2021efficiently, and minimal RNNs, feng2024were. Building on prior work, we propose recurrent transformers, combining the efficacy of transformers with the efficiency of recurrent networks, progressively applying a single transformer layer to a fixed-width sliding window across the input sequence. Prior work on recurrent transformers, e.g., dai2019transformer, where recurrence is used to extend context but does not participate in gradient calculations; and hutchins2022block, focuses on improving the long-range performance of multi-layer transformers by making some of the layers recurrent with large block sizes and many additional weights. In contrast, our focus is to find small and efficient networks for a given task, in the form of a single recurrent layer with a scalar parameter that learns to accumulate or forget history. Specifically, our recurrent transformers (a) run in linear time in the sequence length; (b) are memory-efficient and amenable to parallel processing in large batches (c) learn to forget history for language tasks, or accumulate history for long-range tasks like copy and selective copy; and (d) are amenable to curriculum training to overcome vanishing gradients, e.g.,hochreiter1991untersuchungen, sodhani2020toward. In our experiments, we find that a single-layer recurrent transformer can match the performance of multi-layer transformers on benchmark tests, delivering comparable accuracy with fewer parameters and much lower computational cost. Furthermore, the trainable accumulation capability enables a recurrent transformer to solve long range tasks at a fraction of the size of other recurrent architectures, e.g., gu2023mamba, ren2024exploring, and at a fraction of the cost of regular transformers.
@@ 2 Background
An LLM is a statistical machine that takes as input a sequence of words of specified maximum length on a specified alphabet, and produces as output a probability distribution on possible next words. Training involves adjusting the model’s parameters while running through a body of text called the training corpus, to minimize the statistical error between the model and the corpus. The error in the model, also known as the loss function, is chosen as a continuous function to enable the use of stochastic gradient descent for adjusting the parameters of the model. The commonly used loss function involves the entropy of an infinite stream of words, as defined by shannon1948mathematical, cover1999elements. Intuitively, entropy is the measure of information in the stream and is a lower bound on the capacity of a channel required to transmit the stream.
@@
Let $S$ be the set of all sequences of length $l$ in the training corpus $T$. The Kullback-Leibler divergence of an LLM is
@@
The Kullback-Leibler divergence is well established to be non-negative. Rearranging, we get
@@
The term on the left above is the cross-entropy loss of the LLM on the training corpus. The first term on the right converges to the inherent entropy of the natural language on the training corpus. During the training process, the parameters of the LLM are iteratively adjusted to minimize the cross-entropy. In doing so, the LLM is effectively minimizing the Kullback-Leibler divergence, since the inherent entropy of the language is constant.
@@
All calculations are performed in finite precision such that probabilities $p_{s}$ and $q_{s}$ are quantized over a discrete set of finitely many values $Y\subset[0,1]$, and to avoid overflow in loss calculations the logarithm function is capped, i.e., for all $p,q\in Y$, $|log(p/q)|\leq\lambda$, for some positive constant $\lambda$.
@@ 3 Empirical Scaling
Recent prior works report empirical expressions for the relationship between training loss and model scale, e.g., kmhbccgrwa20hbmbcros22, muennighoff2023scaling.  Per current practice, e.g., hbmbcros22, xfzzy23, these works train LLMs on a single epoch of natural training corpora.  Specifically, hbmbcros22 finds that for fixed sequence length, across a range of LLMs with $N$ parameters and training corpus of size $D$, the Kullback-Leibler divergence scales as follows
@@
In the experiments, the training corpus is randomly split into a training subset and a test subset. The training subset is used for training the LLM, and the Kullback-Liebler divergence is computed on the test subset. Since each training run starts with a random initialization of the neural network, we assume Equation (3) holds with high confidence $(1-\delta)$, for small positive $\delta$ . The first term on the right reflects the ability of an LLM with $N$ parameters to capture the richness of the training corpus. The second term reflects the ability of the training corpus to capture the richness of the natural language.    Since $\alpha\approx\beta,A\approx B$, and the compute cost is $N\times D$, it follows that the optimal scaling strategy is $N\propto D$, which is also known as the AI Scaling Law.    However, $N\propto D$ suggests rote learning, raising the question of whether it is a natural requirement or a reflection of the transformer architecture used in the experiments.
@@ 4 Theoretical Scaling
We now consider LLMs in the context of the theoretical model of Probably Approximately Correct (PAC) learning, v84. Recall that $S$ is the set of sequences of length $l$ in $T$. Consider a hypothesis space of functions $F=\{f:S\rightarrow Y\}$, where for $s\in S$, $f(s)$ is a hypothesis $q_{s}$ for the probabilities $p_{s}$ of Equation 2. We follow sb14 in defining an agnostic learning algorithm as below. As in Section 3, the training corpus $T$ provides $D$ sequences of length $l$ as training samples for the learning algorithm. We define the loss function of the learning algorithm to be the Kullback-Leibler divergence between $f\in F$ and the natural distribution $p_{s}$ as follows
$L(f,T)=\sum_{s\in S}p_{s}log(p_{s}/f(s))$
@@
Given $\epsilon,\delta\in(0,1)$ an agnostic learning algorithm finds $f\in F$ with confidence at least $(1-\delta)$ such that
$L(f,T)\leq\min_{h\in F}L(h,T)+\epsilon$
We assume that the natural distribution follows Zipf’s law, zipf2013psycho for natural languages, which is widely established and valid for sequences, ryland2015zipf.
Assumption 2.
The natural distribution follows Zips’s law in that $p_{s}\propto\frac{1}{r+Z}$, where $r$ is the rank of $s$ in descending order of frequency amongst $s\in S$, and $Z$ is a constant.
The following theorem uses the asymptotic complexity notations $\mathcal{O}$ and $\Omega$ for the upper and lower bound respectively.
@@
For any agnostic learning algorithm for $F$, $\epsilon$ is $\Omega(|S|/D)$ and $\mathcal{O}[(|S|/D)^{0.5}]$
Proof.
We first prove the upper bound. Let $\hat{p}_{s}$ and $\hat{L}(f,T)$ be the observed probabilities and loss respectively in a randomly chosen training subset of the corpus $T$. We have
$|\hat{L}(f,T)|=|\sum_{s\in S}\hat{p}_{s}log(\hat{p}_{s}/f(s))|\leq\sum_{s\in S}|\hat{p}_{s}log(\hat{p}_{s}/f(s))|=\sum_{s\in S}\hat{p}_{s}|log(\hat{p}_{s}/f(s))|$
Invoking Assumption 1 in the above, we get
$|\hat{L}(f,T)|\leq\sum_{s\in S}\hat{p}_{s}|log(\hat{p}_{s}/f(s))|\leq\sum_{s\in S}\hat{p}_{s}\lambda\leq\lambda$
By Hoeffding’s inequality, feller1991introduction, we get
$\mathbb{P}\left\{\left|\hat{L}(f,T)-L(f,T)\right|>\epsilon\right\}\leq 2e^{-2D\epsilon^{2}/\lambda}$
Summing over all choices for $f$, we get
$\sum_{f\in F}\mathbb{P}\left\{\left|\hat{L}(f,T)-L(f,t)\right|>\epsilon\right\}\leq|Y|^{|S|}2e^{-2D\epsilon^{2}/\lambda}$
We wish to bound $\delta$ with the RHS above, i.e., $\delta\leq|Y|^{|S|}2e^{-2D\epsilon^{2}/\lambda}$. Taking logarithms on both sides and treating $\delta,\lambda$, and $|Y|$ as constants yields $\epsilon=\mathcal{O}[(|S|/D])^{0.5}$.
Next we prove the lower bound. For any $f\in F$,
$L(f,T)-\hat{L}(f,T)$ $=\sum_{s\in S}p_{s}log(p_{s}/f(s))-\sum_{s\in S}\hat{p}_{s}log(\hat{p}_{s}/f(s))$
$=\sum_{s\in S}p_{s}log(p_{s}/\hat{p}_{s})-\sum_{s\in S}(\hat{p}_{s}-p_{s})log(\hat{p}_{s})+\sum_{s\in S}(\hat{p}_{s}-p_{s})log(f(s))$
Taking the expectation $\mathbb{E}$ on both sides of the above with respect to random training subsets of T, and noting that $\mathbb{E}\{(\hat{p}_{s}-p_{s})\}=0$, we get,
$\mathbb{E}\left\{L(f,T)-\hat{L}(f,T)\right\}=\mathbb{E}\left\{\sum_{s\in S}p_{s}log(p_{s}/\hat{p}_{s})\right\}$ (4)
Since each sequence $s$ is a Bernoulli variable in a random training subset, the variance of $\hat{p}_{s}$ is $p_{s}(1-p_{s})$. Therefore by the Central Limit Theorem, e.g., feller1991introduction, as $D$ increases without bound, almost surely,
$|\hat{p}_{s}-p_{s}|\propto\sqrt{\frac{p_{s}(1-p_{s})}{D}}$
Invoking Assumption 2 in the above and summing over $s\in S$, we get
$\sum_{s\in S}|\hat{p}_{s}-p_{s}|\propto\frac{1}{\sqrt{D}}\sum_{r=1}^{|S|}\sqrt{\frac{1-\frac{1}{r+Z}}{r+Z}}$
Approximating the sum in the right-hand-side above with the definite integral, we get that
$\sum_{s\in S}|(\hat{p}_{s}-p_{s}|\propto\sqrt{\frac{|S|}{D}}$
Combining the above with Pinsker’s inequality, csiszar2011information, we get
$\sum_{s\in S}p_{s}log(p_{s}/\hat{p}_{s})\geq\frac{1}{2}\left(\sum_{s\in S}|(\hat{p}_{s}-p_{s}|\right)^{2}=\Omega[|S|/D]$
Combining the above with Equation 4, we set
$\zeta=\mathbb{E}\left\{L(f,T)-\hat{L}(f,T)\right\}=\Omega[|S|/D]$
By Assumption 1, $L(f,T)-\hat{L}(f,T)\leq 2\lambda$. Let $x$ be the probability that $L(f,T)-\hat{L}(f,T)>\zeta/2$. The minimum value of $x$ is achieved by placing a probability mass of $x$ at $2\lambda$ and (1x)(1-x) at $\zeta/2$. Hence
$2\lambda x+(\zeta/2)(1-x)=\zeta\implies x(2\lambda-\zeta/2)=\zeta/2\implies x=\frac{\zeta}{4\lambda-\zeta}>\frac{\zeta}{4\lambda}$
It follows that with probability greater than $\frac{\zeta}{4\lambda}$,
$L(f,T)-\hat{L}(f,T)>\zeta/2=\Omega[|S|/D]$
implying that for $f\in F$ output by a learning algorithm, $L(f,T)>\min_{h\in F}L(h,T)+\Omega[|S|/D]$. Hence the theorem.
We now link the empirical bounds of Equation 3 with the theorem above to estimate the number of unique sequences in a training corpus as a function of its size. Our analysis compares the empirical estimates of hbmbcros22 with the estimates of Theorem 1 to extract properties of the underlying natural language.
@@
If Equation 3 holds for sequences of length $l$, the number of unique sequences $|S|$ of length $l$ in a corpus of size $D$ is between $\Omega[D^{0.44}|]$ and $\mathcal{O}[D^{0.72}]$
@@
Consider an LLM with $N$ parameters representing a space of functions $F=\{f:S\rightarrow Y\}$. Suppose the LLM computes a function $f$ after training on corpus $T$.  From Equation 2 and the upper bound of Theorem 1, we have
$\Delta=\sum_{s\in S}p_{s}log(p_{s}/q_{s})=\min_{h\in F}L(h,T)+\mathcal{O}[(|S|/D)^{0.5}]$
@@
$\Delta=\frac{A}{N^{\alpha}}+\frac{B}{D^{\beta}}=\min_{h\in F}L(h,T)+\mathcal{O}[(|S|/D)^{0.5}]$
As the number of parameters $N\rightarrow\infty$, $A/N\rightarrow 0$, and since transformers are universal approximators, $\min_{h\in F}L(h,T)\rightarrow 0$. Hence we have
$\frac{B}{D^{\beta}}=\mathcal{O}[(|S|/D)^{0.5}]$
@@
$|S|=\Omega[D^{1-2\beta}]=\Omega[D^{0.44}].$
Next, we combine Equations 2 with the lower bound of Theorem 1.
$\Delta=\sum_{s\in S}p_{s}log(p_{s}/q_{s})=\min_{h\in F}L(h,T)+\Omega[|S|/D]$
Combining Equations 3 with the above, we get
$\Delta=\frac{A}{N^{\alpha}}+\frac{B}{D^{\beta}}=\min_{h\in F}L(h,T)+\Omega[|S|/D]$
As the number of parameters $N\rightarrow\infty$, $A/N\rightarrow 0$, and since transformers are universal approximators, $\min_{h\in F}L(h,T)\rightarrow 0$. Hence we have
$\frac{B}{D^{\beta}}=\Omega[|S|/D]$
Substituting $\beta=0.28$ from Equation 3 in the above,  we get
$|S|=\mathcal{O}[D^{1-\beta}]=\mathcal{O}[D^{0.72}].$
Which completes the proof.
From a learning theoretic perspective, the number of unique sequences as a function of the size of a training corpus is an upper bound on the degrees of freedom required to learn any natural dataset of that size. Theoretically, the number of parameters in an efficient LLM would scale linearly in the number of unique sequences; e.g., a "brute-force" code book of one entry per unique sequence. The so-called AI scaling law for the transformer architecture suggests that the number of parameters must scale linearly with the size of the data, e.g., kmhbccgrwa20, hbmbcros22, muennighoff2023scaling. In contrast, our main result implies that the number of parameters of an efficient LLM need only scale as $D^{\gamma}$ for $\gamma\in[0.44,0.72]$, suggesting the existence of substantially more efficient LLM architectures.
Our result also has implications for so-called emergent abilities i.e. “abilities that are not present in smaller-scale models but are present in large-scale models; thus they cannot be predicted by simply extrapolating the performance improvements on smaller-scale models” e.g., wei2022emergent, ganguli2022predictability, arora2023theory and srivastava2023beyond. Conversely, schaeffer2023emergent find that LLMs may not exhibit new abilities simply by increasing the number of parameters. Our result implies that if the number of parameters of an LLM is smaller than the number of unique sequences in the training corpus, scaling up the LLM can uncover emergent skills.